OpenStudy (anonymous):

Which fraction is equivalent to 4 2/3? MEDAL + FANNNN

OpenStudy (wyattp17):

4.66 is the decimal

OpenStudy (eli_moses):

14/3

OpenStudy (wyattp17):

so just find what is equal to 4.66

OpenStudy (kendricklamar2014):

14/3

OpenStudy (anonymous):

14/3

OpenStudy (kendricklamar2014):

4 2/3 = 12/3 + 2/3 = 14/3 So 4 2/3 is equivalent to 14/3